# Local Reinforcement Learning with Action-Conditioned Root Mean Squared Q-Functions

### v1

@@ 5 Experiments
We test ARQ on the MinAtar benchmark (Young and Tian,, 2019) and the DeepMind Control (DMC) Suite (Tassa et al.,, 2018) following Guan et al., (2024). MinAtar is a miniaturized version of the Atari 2600 games, using 10x10 grids instead of 210x160 frames as inputs. The DMC Suite is a benchmark for continuous control tasks featuring low-level observations and actions, designed to evaluate the performance of RL methods in physics-based environments. Both benchmarks involve low-dimensional inputs and outputs instead of high-dimensional raw sensory inputs, making them appropriate testbeds for evaluating the decision-making ability of local methods in simple environments.
@@
For comparisons with cutting-edge local RL methods, we compare our results with AD for both benchmarks. To evaluate our methods against backprop-based algorithms, we also compare our method against DQN (Mnih et al.,, 2015) for MinAtar. DQN is a widely used baseline that trains deep neural networks to directly compute scalar Q-values through backpropagation. We follow the DQN implementation used by Guan et al., (2024).
Figure 4: Training performance on the MinAtar games, compared between DQN (blue), AD (orange), and ARQ(green). The x-axis denotes the number of training steps (in millions), and the y-axis indicates average episodic returns. Shaded regions represent standard deviations across 3 seeds. We find that ARQ consistently outperforms AD in all MinAtar games, while outperforming DQN in most games.
@@
Following Guan et al., (2024), we use a three-layer fully-connected network, with hidden dimensions being 400, 200, and 200 for MinAtar. We use a three-layer network with hidden dimensions 128, 96, and 96 for DMC tasks. We use a replay buffer and a target network for stability. We incorporate skip connections from the input and top-down connections from the layer above. For all experiments, we use an epsilon-greedy policy with linear decay from 1 to 0.01 using an exploration fraction of 0.1. We run our experiments for 4 million steps, where the model starts learning from step 50,000. Learning rate is set fixed at 1e-4. A batch size of 512 is used. For MinAtar, we condition on action candidates by passing them as one-hot vectors into the network. For DMC tasks, we discretize our action space and condition action vectors as model inputs.
@@
We present our results in Table 1. We run each experiment with three different random seeds and plot their average returns over 100-episode windows along with their standard deviations in shadows. We also calculated the average returns of the last 100 episodes of each training run to obtain a quantitative measure of the final performance of our method, which can be found in Table 1. As demonstrated, ARQ consistently outperforms AD in all MinAtar games. Surprisingly, ARQ also outperforms DQN in all games. In DMC Suite tasks, ARQ achieves superior returns compared to AD, while also exceeding back-prop based methods in most games. We present our numbers in Table 1.
@@
Table 1: Previous methods and ARQ compared in MinAtar and DeepMind Control (DMC) tasks.
@@
Table 2: Our method Using Different Nonlinearities Compared in MinAtar Breakout. ‘MS’ is short for the mean squared function and ‘Var’ is short for variance. Default ARQ uses the root mean squared (RMS) function.
@@
AD 67.40 $\pm$ 4.01 369.96 $\pm$ 23.46
Effect of Action Conditioning at Input:
How does action conditioning affect the performance of local RL methods? To investigate, we conduct ablation experiments on two games from MinAtar, Breakout and SpaceInvaders, using both AD and ARQ. The results can be found in Figure 5. We find a significant improvement when actions are conditioned at the input instead of at the output. Interestingly, this design choice provides only a slight improvement for AD, while yielding a significant increase in performance for ARQ. We conjecture this is due to the increase in the capacity of each cell to capture the granularity within each specific state-action pair, while AD saturates with action-agnostic information.
@@
One question that naturally arises is the choice of the goodness function. Does the RMS function perform superiorly compared to other functions? We ablate on this design choice and conduct experiments on two games from MinAtar, Breakout and SpaceInvaders. As shown in Table 2, we find that using the RMS goodness functions yields superior performance, followed by the mean and the mean squared function. We conjecture that a smaller magnitude in the goodness can enhance stability of training. However, we note that all functions perform superiorly compared with AD, which demonstrates the versatility of our method. We leave it for future work to study the intricate effect each function has on training.
Table 3: AD vs. ARQ Across Multiple Scales for MinAtar Breakout.
@@
Is it because ARQ has more hidden units?
Figure 6: Visualization of Neurons in Layer 0 under Different Scenarios in Breakout Game. 20 neurons w/ highest average activities are visualized. Top Left: When the ball is approaching towards the left side of the brick, neurons show larger magnitude when the action candidate is to “move left”, prompting the agent to move towards the ball. Bottom Left: When the ball approaches the right side of the brick, neurons show larger magnitude when the action candidate is to “move right”. Right: The average root mean squared (RMS) activations of 20 top neurons across 100 states is collected. Note that top neurons exhibit significantly larger RMS activations than the average RMS activation, implying that these neurons are “dominant” neurons. While most neurons demonstrate similar magnitude between both actions, some neurons appear to be more specialized.
@@
How does our method learn through a goodness function? We investigate its inner mechanism by visualizing the activations at each layer under different states. As illustrated in Figure 6, we find that the hidden activations tend to show larger magnitudes under “correct” state-action pairs. For instance, in scenarios where the agent should move right to accurately catch the incoming ball, neurons in the hidden activations show the largest magnitude when the action input matches correspondingly. Interestingly, we observe that different neurons are, in general, activated to different degrees for various action candidates. This implies our objective function could be encouraging specialized neurons, each of which is responsible for recognizing certain categories of positive signals.
@@ 6 Discussion
Previous studies on biologically plausible learning have largely focused on the search for a biologically plausible mechanism for performing gradient updates. As we approach the era of AIs with agentic learning and experience, we argue that a biologically plausible learning environment can be equally meaningful in guiding us towards understanding the mystery behind how biological brains learn. Reward-centric environments provide a biologically grounded environment that aligns with the evolutionary role of survival signals and behavioral shaping through positive or negative reinforcement. The structure of such environments mirrors the ecological settings in which animals adaptively refine behavior through trial-and-error interactions, suggesting that learning systems shaped by rewards may naturally emerge in both artificial and biological agents. Additionally, temporal difference methods are ideal candidates as there exist evidence showing that biological neurons learn through temporal difference, with hormones conveying the prediction error as a source of the learning signal to independent neurons (Schultz et al., 1997b, ; Bayer and Glimcher,, 2005). On the other hand, reinforcement learning has largely focused on learning through interactions with an agent’s surrounding environment and maximizing its rewards through centralized value estimation. Yet, increasing neuroscientific evidence has shown that neurons make decentralized, independent value estimations (Tsutsui et al.,, 2016; Knutson et al.,, 2005). Few studies in the RL community have investigated whether this biological phenomenon has practical implications. ARQ is an effort towards this direction, as each cell in our network can be seen as a decentralized value estimator.
@@ 7 Conclusion
This work proposes Action-conditioned Root mean squared Q-function (ARQ), a vector-based alternative to scalar Q-learning for backprop-free local learning. ARQ enables arbitrary hidden dimensions and improved expressivity by extracting value predictions from hidden activations and applying action conditioning at the model input. We show that, when applied on RL environments, ARQ performs superiorly compared to current local methods, while also outperforming backprop-based methods on most games. Whereas current biologically plausible algorithms are mostly based on the supervised setting, our study suggests that exploring local learning within reinforcement learning may provide a promising avenue for future research in both domains.

### v2

@@ 5 Experiments
We test ARQ on the MinAtar benchmark (Young and Tian, 2019) and the DeepMind Control (DMC) Suite (Tassa et al., 2018) following Guan et al. (2024). MinAtar is a miniaturized version of the Atari 2600 games, using 10x10 grids instead of 210x160 frames as inputs. The DMC Suite is a benchmark for continuous control tasks featuring low-level observations and actions, designed to evaluate the performance of RL methods in physics-based environments. Both benchmarks involve low-dimensional inputs and outputs instead of high-dimensional raw sensory inputs, making them appropriate testbeds for evaluating the decision-making ability of local methods in simple environments.
@@
For comparisons with cutting-edge local RL methods, we compare our results with AD for both benchmarks. To evaluate our methods against backprop-based algorithms, we also compare our method against DQN for MinAtar. DQN is a widely used baseline that trains deep neural networks to directly compute scalar Q-values through backpropagation. We follow the DQN implementation used by Guan et al. (2024).
Figure 4: Training performance on the MinAtar games, compared between DQN (blue), AD (red), and ARQ (green). The x-axis denotes the number of training steps (in millions), and the y-axis indicates average episodic returns. Shaded regions represent 95% confidence intervals across 5 seeds. We find that ARQ consistently outperforms AD in all MinAtar games, while outperforming DQN in some games.
@@
Following Guan et al. (2024), we use a three-layer fully-connected network, with hidden dimensions being 400, 200, and 200 for MinAtar. We use a three-layer network with hidden dimensions 128, 96, and 96 for DMC tasks. We use a replay buffer and a target network for stability. We incorporate skip connections from the input and top-down connections from the layer above. For all experiments, we use an epsilon-greedy policy with linear decay from 1 to 0.01 using an exploration fraction of 0.1. We run our experiments for 4 million steps, where the model starts learning from step 50,000. Learning rate is set fixed at 1e-4. A batch size of 512 is used. For MinAtar, we condition on action candidates by passing them as one-hot vectors into the network. For DMC tasks, we discretize our action space and condition action vectors as model inputs.
@@
As presented in Figure 4, we run each experiment with five different random seeds and plot their average returns over 100-episode windows along with their 95% confidence intervals in shadows. We also calculated the average returns of the last 100 episodes of each training run to obtain a quantitative measure of the final performance of our method, which can be found in Table 1. As demonstrated, ARQ consistently outperforms AD in all MinAtar games. Surprisingly, ARQ also outperforms DQN in all games. In DMC Suite tasks, ARQ achieves superior returns compared to AD, while also exceeding back-prop based methods in most games. A possible explanation is that ARQ benefits from localized TD updates, reduced gradient path length, and the variance reduction effect of layerwise averaging, which together can lead to more stable and efficient learning than fully backpropagated networks.
@@
Table 1: Performance of previous methods and ARQ on MinAtar and DeepMind Control (DMC) tasks. Reported as mean ± 95% confidence intervals across 5 random seeds.
@@
Effect of Action Conditioning at Input:
How does action conditioning affect the performance of local RL methods? To investigate, we conduct ablation experiments on two games from MinAtar, Breakout and SpaceInvaders, using both AD and ARQ. The results can be found in Figure 5. We find a significant improvement when actions are conditioned at the input instead of at the output. To further understand this difference, we analyzed the hidden activations using PCA and compared them against predicted Q-values as presented in Figure 6. Without action conditioning, activations cluster almost entirely by action identity and show no meaningful correlation with Q-values, indicating that action-related variance dominates the representation space. With action conditioning, representations become more state-driven and exhibit a mild positive relationship with Q-values, suggesting that the model can allocate capacity toward value-relevant structure rather than implicitly inferring action identity. Interestingly, this design choice provides only a slight improvement for AD, while yielding a significant increase in performance for ARQ. We conjecture this is due to the increase in the capacity of each cell to capture the granularity within each specific state–action pair, while AD saturates with action-agnostic information.
Figure 6: Representation analysis of ARQ with and without action conditioning. We train two agents—ARQ with and without action conditioning (AC)—on MinAtar Breakout, then randomly sample 200 states from each trained policy. For each state–action pair, we extract the hidden activations from Layer 0 and visualize them using 2-component PCA. The top row shows how mPCA components cluster across different actions. Without AC, activations form tight clusters determined almost entirely by action identity, indicating that the agent must implicitly encode action information inside the representation. With AC, the activations are more entangled and state-driven. The bottom row plots the first PCA component against the predicted Q-values. With AC, ARQ exhibits a mild positive correlation between latent structure and Q-values, while the non-AC model shows no meaningful correlation and remains dominated by action-specific clustering.
Table 2: Our method Using Different Nonlinearities Compared in MinAtar Breakout. ‘MS’ is short for the mean squared function and ‘Var’ is short for variance. Default ARQ uses the root mean squared (RMS) function. Reported as mean ± 95% confidence intervals.
@@
AD 67.40 $\pm$ 8.02 369.96 $\pm$ 46.92
@@
One question that naturally arises is the choice of the goodness function. Does the RMS function perform superiorly compared to other functions? We ablate on this design choice and conduct experiments on two games from MinAtar, Breakout and SpaceInvaders. As shown in Table 2, we find that using the RMS goodness functions yields superior performance, followed by the mean and the mean squared function. Beyond performance, our analysis suggests that RMS maintains healthier activation magnitudes throughout training, whereas using mean squared function produces extremely large early goodness values that later suppress activation norms (see Figure 7). This stabilization effect likely preserves a richer and more expressive representation space, contributing to RMS’s empirical advantage. However, we note that all functions perform superiorly compared with AD, which demonstrates the versatility of our method.
Figure 7: Analysis of ARQ (RMS goodness) vs. ARQ-MS (mean-squared goodness) on MinAtar–Breakout. We evaluate hidden activations and goodness values of both agents over 200 randomly sampled states at both the start and end of training. Top row: Distribution of L2 norms of hidden activations. Early in training, ARQ-MS exhibits extremely large goodness values and compressed activation magnitudes, while ARQ maintains broader, more expressive activation scales. By the end of training, ARQ continues to support significantly larger activation norms, whereas ARQ-MS activations remain narrowly concentrated. Bottom row: Distribution of scalar goodness values derived from each hidden vector. ARQ produces moderate, stable goodness magnitudes throughout training, while ARQ-MS shows large initial spikes followed by sharply reduced variability.
Is it because ARQ has more hidden units?
Table 3: AD vs. ARQ Across Multiple Scales for MinAtar Breakout. Reported as mean ± 95% confidence intervals.
@@
How does our method learn through a goodness function? We investigate its inner mechanism by visualizing the activations at each layer under different states. As illustrated in Figure 8, we find that the hidden activations tend to show larger magnitudes under ”correct” state-action pairs. For instance, in scenarios where the agent should move right to accurately catch the incoming ball, neurons in the hidden activations show the largest magnitude when the action input matches correspondingly. Interestingly, we observe that different neurons are, in general, activated to different degrees for various action candidates. This implies our objective function could be encouraging specialized neurons, each of which is responsible for recognizing certain categories of positive signals.
@@ 6 Discussion
Previous studies on biologically plausible learning have largely focused on the search for a biologically plausible mechanism for performing gradient updates. As we approach the era of experience, we argue that a biologically plausible paradigm for learning can be equally meaningful to guide us towards the mystery behind how biological brains learn. Reward-centric environments provide a biologically grounded paradigm, aligning with the evolutionary role of survival signals and behavioral shaping through positive or negative reinforcement. The structure of such environments mirrors the ecological settings in which animals adaptively refine behavior through trial-and-error interactions, suggesting that learning systems shaped by rewards may naturally emerge in both artificial and biological agents. Additionally, temporal difference methods are an ideal candidate. It has been shown that biological neurons learn through temporal difference, with hormones conveying the prediction error as a source of learning signal to independent neurons (Schultz et al., 1997b; Bayer and Glimcher, 2005). On the other hand, reinforcement learning has largely focused on learning through interactions with an agent’s surrounding environment, and maximizing its rewards through centralized value estimation. Yet, increasing neuroscientific evidence has shown that neurons make decentralized, independent value estimations (Tsutsui et al., 2016; Knutson et al., 2005). Few work in the RL community has investigated whether this biological phenomenon has practical implications. ARQ is an effort towards this direction as each cell in our network can be seen as a decentralized value estimator.
@@ 7 Conclusion
This work proposes Action-conditioned Root mean squared Q-Function (ARQ), a vector-based alternative to scalar Q-learning for backprop-free local learning. ARQ enables arbitrary hidden dimensions and improved expressivity by extracting value predictions from hidden activations and applying action conditioning at the model input. We show that, when applied on RL environments, ARQ performs superiorly compared to current local methods, while also outperforming backprop-based methods on some games. Whereas current biologically plausible algorithms are mostly based on the supervised setting, our study suggests that exploring local learning within reinforcement learning may provide a promising avenue for future research in both domains.
@@ Acknowledgement
Appendix A Visualization of ARQ Activations
Figure 8 provides the full activation visualizations referenced in the main text, illustrating how neuron responses vary across different state–action scenarios in Breakout.
Figure 8: Visualization of Neurons in Layer 0 under Different Scenarios in Breakout Game. 20 neurons w/ highest average activities are visualized. Top Left: When the ball is approaching towards the left side of the brick, neurons show larger magnitude when the action candidate is to “move left”, prompting the agent to move towards the ball. Bottom Left: When the ball approaches the right side of the brick, neurons show larger magnitude when the action candidate is to ”move right”. Right: The average root mean squared (RMS) activations of 20 top neurons across 100 states is collected. Note that top neurons exhibit significantly larger RMS activations than the average RMS activation, implying that these neurons are ”dominant” neurons. While most neurons demonstrate similar magnitude between both actions, some neurons appear to be more specialized.
Appendix B Hyperparameters
We summarize all hyperparameters used for MinAtar and DeepMind Control Suite experiments in Table 4, consolidating the network architectures, training settings, and optimization details for full reproducibility.
Hyperparameter MinAtar DMC Suite
Network Architecture 3-layer MLP 3-layer MLP
Hidden Dimensions 400-200-200 128-96-96
Optimizer Adam Adam
Discount Factor ($\gamma$) 0.99 0.99
Learning Rate $\times 10^{-4} 1×1$$\times 10^{-4}$
Batch Size 512 512
Replay Buffer Size 4M transitions 4M transitions
Target Network Yes Yes
Exploration Strategy $\epsilon$-greedy $\epsilon$-greedy (on discretized actions)
$\epsilon$ Schedule Linear: 1.0 $\rightarrow$ 0.05 Linear: 1.0 $\rightarrow$ 0.05
Exploration Fraction 0.1 0.1
Learning Starts 50,000 steps 50,000 steps
Training Steps 4M 4M
Action Representation One-hot Bang-bang discretization
Random Seeds 5 5
Table 4: Consolidated hyperparameters for MinAtar and DeepMind Control (DMC) experiments. All settings follow the implementation described in Section 5 of the paper, with Adam optimization, a shared replay buffer of 4M transitions, and identical training schedules. Action conditioning is applied for ARQ by default.
Appendix C Detailed Computational Diagram
In this section, we provide a detailed visualization of the computation flow in both AD and ARQ, illustrating how information propagates across layers and how each cell locally estimates Q-values.
Figure 9: Detailed computation diagram. Key implementations of ARQ are highlighted in red. Left: Overall information processing across layers. Each cell receives input from the raw observation, the layer below, and the layer above, and its activations are passed forward to the next temporal step. Top Right: Cell Computation in Artificial Dopamine (AD); $n_{a}$ Q-value estimates are produced by dot-products between the attentional weights and the hidden states. Bottom Right: Our ARQ implements root mean squared functions for value estimation along with action-conditioned inputs.
Appendix D DeepMind Control Suite Experimental Figures
Figure 10 presents the full training curves for all DeepMind Control Suite environments. ARQ consistently outperforms AD across tasks and achieves performance comparable to strong backpropagation-based methods such as TD-MPC2 and SAC.
Figure 10: Training performance on the DeepMind Control Suite, compared between AD (red), ARQ (green), TD-MPC2 (blue), and SAC (gray). The x-axis shows training steps (in millions), and the y-axis denotes average episodic returns. Shaded regions indicate 95% confidence intervals across 5 seeds. Across all environments, ARQ consistently improves over AD and achieves performance competitive with backpropagation-based methods.
Appendix E Nonlinearity Ablation Figures
Figure 11 presents the per-environment learning curves comparing different nonlinear goodness functions used to extract scalar values from hidden vectors. The default ARQ, which uses the root mean squared function, outperforms all other choices. Notably, all variants outperform AD in both settings.
Figure 11: Comparison of different nonlinear goodness functions that may be used to collect scalar statistics from hidden vectors. ‘MS’ is short for the mean squared function and ‘Var’ is short for variance. Default ARQ uses the root mean squared (RMS) function. Shaded regions represent 95% confidence intervals.